Discount-Mart, a major East Coast retailer, wants to determine the economic order quantity for its halogen lamps. It currently buys all halogen lamps from Specialty Lighting Manufacturers, in Atlanta. Annual demand is 2,000 lamps, ordering cost per order is $60, carrying cost per lamp is $12 a) What is the EOQ? 142 lamps per order (round your response to the nearest whole number). b) What are the total annual costs of holding and ordering (managing) this inventory? $ 1,697 (round your response to the nearest whole number). c) How many orders should Discount-Mart place with Specialty Lighting per year? ? orders per year round your response to the nearest whole number
Answer #1

Demand (D)=2,000 units

Holding Cost (H) = $12 per unit per year

Setup cost (S)=$60

a) EOQ= (2DS/H)^0.5 = (2*2000*60/12)^0.5= 141.42= 142 lamps

b) Total annual cost = SD/Q + HQ/2 = (60*2000/142) + (12*142/2) = $1697.07= $1697

c) No of orders= D/Q = 2000/142 = 14.08 = 15 orders per year